# Reinforcement Learning vs. Distillation: Understanding Accuracy and Capability in LLM Reasoning

### v1

@@ Abstract
Recent studies have shown that reinforcement learning with verifiable rewards (RLVR) enhances overall accuracy but fails to improve capability, while distillation can improve both. In this paper, we investigate the mechanisms behind these phenomena. First, we demonstrate that RLVR does not improve capability because it focuses on improving the accuracy of the less-difficult questions to the detriment of the accuracy of the most difficult questions, thereby leading to no improvement in capability. Second, we find that RLVR does not merely increase the success probability for the less difficult questions, but in our small model settings produces quality responses that were absent in its output distribution before training. In addition, we show these responses are neither noticeably longer nor feature more reflection-related keywords, underscoring the need for more reliable indicators of response quality. Third, we show that while distillation reliably improves accuracy by learning strong reasoning patterns, it only improves capability when new knowledge is introduced. Moreover, when distilling only with reasoning patterns and no new knowledge, the accuracy of the less-difficult questions improves to the detriment of the most difficult questions, similar to RLVR. Together, these findings offer a clearer understanding of how RLVR and distillation shape reasoning behavior in language models. [^1] Code available at https://github.com/minwukim/RLvsDistillation
@@
Minwu Kim*  Anubhav Shrestha*  Safal Shrestha  Aadim Nepal  Keith Ross New York University Abu Dhabi
11footnotetext: Equal contribution.footnotetext: Correspondence: keithwross@nyu.edu
@@ 1 Introduction
A central technique behind recent reasoning models is reinforcement learning with verifiable rewards (RLVR) Lambert et al. (2024); DeepSeek-AI (2025). RLVR fine-tunes pre-trained LLMs using rewards based on whether the model’s output matches a ground-truth solution. Without explicit supervision, the model learns complex reasoning behaviors during training, making RLVR an effective fine-tuning method for reasoning tasks. In addition, distillation with responses from stronger reasoning models also provide strong performance, demonstrating that reasoning ability can be effectively transferred through supervised learning Min et al. (2024); DeepSeek-AI (2025); Muennighoff et al. (2025); Ye et al. (2025).
It has been recently observed that RLVR improves accuracy—the probability that a correct answer is generated, but does not improve capability-whether the correct answers exists within the model’s output distribution Yue et al. (2025); Shao et al. (2024). It has been further observed that distillation improves both accuracy and capability Yue et al. (2025). This prior research points out important “whats” for RLVR and teacher distillation, and provides an interesting contrast among the two techniques. In this paper, we take a deeper dive into these issues by examining not only the “whats” but also “why” these phenomena occur.
@@
Carrying out experiments with two models, Qwen2.5-1.5B-Math Yang et al. (2024) and Qwen2.5-3B Hui et al. (2024), we first demonstrate that RLVR does not improve capability because RLVR focuses on improving the accuracy of the less-difficult questions to the detriment of the accuracy of the most difficult questions, explaining why capability does not improve and can even decrease. We argue that this "sacrificing-difficult-problems" phenomenon is a direct consequence of the underlying policy-gradient algorithm in GRPO Shao et al. (2024). We further find that RLVR does not merely increase the success probability for the less difficult questions, but in our small model settings produces responses that are more direct with fewer keywords. In addition, we find these responses are neither noticeably longer nor richer in reflection-related keywords (e.g., “wait”, “alternatively”), underscoring the need for more reliable indicators of reasoning quality.
Finally we turn our attention to teacher distillation. A teacher model’s responses carry two main types of information: (1) its reasoning patterns and (2) its domain knowledge. To disentangle their effects, we compare three models: the base model; the publicly released DeepSeek reasoning model, which was reportedly distilled 800k responses from DeepSeek-R1 and likely incorporates substantial new knowledge; and our own distilled model, trained only on teacher responses for questions whose correct answers already existed in the base model’s output distribution. We find that although both distilled models significantly improve accuracy, only the DeepSeek model shows a meaningful increase in capability. Thus the reason why teacher distillation improves both accuracy and capability is not because it brings new reasoning patterns to the distilled model but instead brings new knowledge, knowledge which was not present in the original model. Comparing RLVR and teacher distillation is thus an apples-to-oranges comparison, in that RLVR only uses the resources it has on hand, whereas teacher distillation augments the model with new resources. We also make one other interesting observation: when distilling only with reasoning patterns and no new knowledge, the accuracy of the less-difficult questions improves to the detriment of the most difficult questions, similar to RLVR.
Taken together, our findings provide a clearer picture of behind the "whys" of RLVR-trained and teacher-distilled reasoning models and help to contrast the two approaches.
@@ 2 Related Work
2.1 Reinforcement Learning with Verifiable Rewards (RLVR)
RLVR has emerged as an effective method for training LLMs to solve complex reasoning tasks Lambert et al. (2024). Notable demonstrations include OpenAI’s o1 OpenAI (2024), DeepSeek-R1 DeepSeek-AI (2025), and Kimi k1.5 MoonshotAI (2025), which use RLVR to produce long CoTs and achieve substantial gains across various benchmarks. Follow-up studies have replicated these successes across different model sizes Gandhi et al. (2025); Hu et al. (2025); Liu et al. (2025); Xu et al. (2025); Yeo et al. (2025); Zeng et al. (2025), as well as in diverse domains such as medical and synthetic reasoning tasks Pan et al. (2025); Xie et al. (2025); Zhang et al. (2025). In parallel, various algorithmic variants of RLVR have been proposed to improve model performance, data efficiency, and computational cost Shao et al. (2024); Fatemi et al. (2025); Liu et al. (2025); Wang et al. (2025); Yu et al. (2025); Zuo et al. (2025). However, most prior work has focused on improving the effectiveness of RLVR itself. In contrast, our work aims to deepen the understanding of how RLVR changes model behavior. By analyzing its effects through the dual perspectives of accuracy and capability, we clarify the mechanisms by which RLVR shapes reasoning in LLMs.
2.2 Distillation for Reasoning in LLMs
Distilling high-quality chain-of-thought (CoT) data has also proven to be an effective approach for improving the reasoning ability of LLMs. Various efforts have focused on collecting such data. In some cases, high-quality reasoning traces are generated by prompting large language models Zelikman et al. (2022); Yu et al. (2024); in others, human annotators construct answers that reflect complex, non-linear reasoning patterns Ye et al. (2025); Qin et al. (2024); Xiang et al. (2025). More recently, a popular strategy involves distilling responses from RLVR-trained reasoning models into student models, often resulting in significant performance improvements Huang et al. (2024); Min et al. (2024); Muennighoff et al. (2025). In parallel, several studies have explored which qualitative properties of CoT demonstrations—such as problem difficulty or structural format—most strongly influence the effectiveness of distillation Li et al. (2025); Sun et al. (2025). What sets our work apart is a fine-grained analysis of the distillation data itself. We disentangle the effects of reasoning style and domain knowledge, examining how each influences both model accuracy and capability. This perspective provides deeper insight into how reasoning-focused distillation shapes model behavior.
@@ 3.1 Formal Definitions
We evaluate models along two dimensions: accuracy and capability. Informally, accuracy measures how likely a model is to generate a correct answer in a single attempt, while capability measures whether a correct answer exists within the distribution of responses the model can generate.
@@ 4 Why Doesn’t RLVR Improve Capability?
In prior work, it was shown empirically from various models that RLVR can provide significant improvements in accuracy but in doing so it may not improve pass@$k$ capability when $k$ is chosen to be sufficiently large Yue et al. (2025). In this section, we build on that observation and examine the phenomenon in greater depth. Our analysis also lays the foundation for the experiments that follow. Specifically, we aim to answer the following question: Why does RLVR improve accuracy while leaving capability unchanged or even degraded?
@@ 4.2 A Deeper-Dive: Analysis Based on Question Difficulty
To better understand the accuracy-capability dynamics of RLVR, we perform a fine-grained analysis based on question-level success rates. Specifically, we perform the following experiment: using the 256 responses generated by the base model for each question in the training and test sets, we compute a per-question success rate. We then group the questions into bins based on these success rates: [0], [1–4], [5–16], [17–64], [65–128], and [129-256]. For each bin, we identify the corresponding questions and retrieve the RL model’s responses to those same questions. Finally, we compute the average success rate for both the base and RL models within each bin, and plot the difference between them. The results are shown in Figure 1.
@@
We observe two consistent patterns. (1) For questions where the base model already has a moderately high success rate, the RL model shows substantial improvement. For example, in the training set, the [65–128] bin shows a 36.6% gain in average success rate relative to 256 responses, and the [17–64] bin shows a 24.7% gain. (2) In contrast, for questions where the base model has zero or near-zero success rate, the improvement is negligible. For instance, the [0] bin in the training set improves by only 0.5% points, and the [1–4] bin by just 0.6% points. These patterns are consistent across both the train and test sets and across both model sizes (1.5B and 3B), as shown in Appendix A.3.
@@
Notably, we observe two clear trends. (1) Questions that were already in high-success bins tend to stay in the same bin or move to even higher bins after RLVR training. For instance, in the [65–128] bin, 15.4% remain in place and striking 83.3% move to higher [129-256] bin—while only 1.3% (1 question out of 78) drop to lower ones. Similar upward shifts are seen in the [17–64] bin. (2) In contrast, questions that started in low-success bins, especially those with near-zero success rates, are more likely to stay in the same bin or shift further downward. For example, in the [1-4] bin, a substantial 61.1% remain in the same bin while 16.7% drop down to the [0] bin, and a similar pattern appears in the [5–16] range where 67.4% of the questions stay in the same bin or regress. This shows a clearer picture of RLVR failing to help with previously unsolved questions and can even increase their number. Many questions with a small chance of being answered correctly end up never being answered correctly at all after training. Again, these trends are consistent across both the train and test sets and across both 1.5B and 3B models, as discussed in Appendix A.3.
To understand this behavior, we need to examine the internal dynamics of RLVR training. Take GRPO algorithm as an example. At each iteration, the model generates multiple responses (e.g., 8) per question in a batch. Each response is evaluated for correctness, and parameter updates are applied accordingly. If all 8 responses for a given question are incorrect, that question has no influence on the update. In contrast, for questions with a mix of correct and incorrect responses, the model parameters are nudged to increase the likelihood of generating correct answers and decrease incorrect ones. Now consider an early training step where the batch includes one difficult question (e.g., in bin [1–4]) and several easier ones (e.g., in bin [65–128]). With high probability, the model will generate only incorrect responses for the difficult question, while producing at least some correct answers for the easier ones. As a result, the parameter update will be guided entirely by the easier questions only. This can lead to the model assigning even lower probability to the correct answers for the difficult question, especially if model capacity is limited. This selective reinforcement continues throughout training, explaining why questions that initially had a small chance of being answered correctly may become even less likely to be solved after training. To confirm this selective reinforcement effect, we also analyze the entropy of the model’s output distributions—measured over 256 sampled responses per question—across different difficulty levels. We observe that entropy consistently decreases after RLVR training. Full results are provided in Appendix A.4.
To summarize, these results suggest to the following insight: RLVR improves accuracy but not capability because RLVR focuses on improving the accuracy of the less-difficult questions to the detriment of the accuracy of the most difficult questions.
@@ 5.1 Self-Distillation Experiments
Table 1: Self-distillation results for the Qwen2.5-1.5B-Math base model and its RLVR-trained variant. The notation $A$leftarrow $B$ indicates the student model $A$ is trained on responses from the teacher model $B$. Values in parentheses show gains over the student model with distillation.
@@
We take a step further and conduct another self-distillation experiment. Prior work has shown that distilling quality responses from a teacher model can effectively improve the performance of student model  Huang et al. (2024); Min et al. (2024); Muennighoff et al. (2025). In other words, responses that lead to large accuracy gains through distillation can be seen as having high quality. Based on this idea, we perform an additional experiment: distilling the RL model’s responses into the base model.
@@
In summary, when distilling the base model with correct responses from the base model there is only minor improvement, whereas when distilling with correct responses from the RLVR-trained model there is significant improvement, on par with RLVR itself. This suggests that there is a qualitative difference in the two response types, and reveal the following insight: RLVR does more than merely increase the success probability for the less difficult questions—it enables the model to produce quality responses that were not present in its output distribution before training.

### v2

@@ Abstract
Recent studies have shown that reinforcement learning with verifiable rewards (RLVR) enhances overall accuracy (pass@1) but often fails to improve capability (pass@$k$) of LLMs in reasoning tasks, while distillation can improve both. In this paper, we investigate the mechanisms behind these phenomena. First, we demonstrate that RLVR struggles to improve capability as it focuses on improving the accuracy of the easier questions to the detriment of the accuracy of the most difficult questions. Second, we show that RLVR does not merely increase the success probability for the easier questions, but in our small model settings, produces quality responses that were absent in its original output distribution. In addition, we show these responses are neither noticeably longer nor feature more reflection-related keywords, underscoring the need for more reliable indicators of response quality. Third, from the experiment distilling teacher responses to in-distribution problems, we find that capability does not always improve with distillation. We conjecture that capability improves only when new knowledge is introduced, whereas distilling reasoning patterns only improves accuracy but not capability, sacrificing performance on the most difficult questions, similar to RLVR. Together, these findings offer a clearer understanding of how RLVR and distillation shape reasoning behavior in LLMs. [^1] Code available at https://github.com/minwukim/RLvsDistillation
@@
Minwu Kimthanks: Equal contribution. thanks: Correspondence to mwk300@nyu.edu.   Anubhav Shrestha*   Safal Shrestha   Aadim Nepal   Keith Ross New York University Abu Dhabi
@@ 1 Introduction
A central technique behind recent reasoning models is reinforcement learning with verifiable rewards (RLVR) Lambert et al. (2024); DeepSeek-AI (2025). RLVR fine-tunes pre-trained LLMs using rewards based on whether the model’s output matches a ground-truth solution. Without explicit supervision, the model learns complex reasoning behaviors during training, making RLVR an effective fine-tuning method for reasoning tasks. In addition, distillation with responses from stronger reasoning models also provides strong performance, demonstrating that reasoning ability can be effectively transferred through supervised learning Min et al. (2024); DeepSeek-AI (2025); Muennighoff et al. (2025); Ye et al. (2025).
It is well established that RLVR improves accuracy—the probability of generating a correct answer, but whether it also improves capability—the probability that a correct answer exists in the model’s output distribution—remains debated. Some studies suggest that, with sufficient compute and carefully matched training and test sets in skills and difficulty, RLVR can solve tasks that were previously unsolvable in certain domains Liu et al. (2025a); Setlur et al. (2025); Sun et al. (2025). Others, however, report that in more typical settings—where training and test sets contain heterogeneous problems with uncontrolled knowledge and difficulty—RLVR primarily amplifies existing reasoning rather than expanding capability Dang et al. (2025); Wu et al. (2025); Yue et al. (2025); Zhao et al. (2025). By contrast, it has been observed that distillation improves both accuracy and capability Yue et al. (2025). In this paper, we take a closer look at how RLVR and distillation shape mathematical reasoning in LLMs under typical settings, where training and test sets involve diverse problems with varying knowledge and difficulty.
@@
Carrying out experiments with two models, Qwen2.5-1.5B-Math Yang et al. (2024) and Qwen2.5-3B Hui et al. (2024), we first demonstrate that RLVR does not improve capability because RLVR focuses on improving the accuracy of the less-difficult questions to the detriment of the accuracy of the most difficult questions, explaining why capability does not improve and can even decrease. We argue that this "sacrificing-difficult-problems" phenomenon is a direct consequence of the underlying policy-gradient algorithm in GRPO (Shao et al., 2024). We further find that RLVR does not merely increase the success probability for the easier questions, but in our small model settings produces responses that are more direct with fewer keywords. In addition, we find these responses are neither noticeably longer nor richer in reflection-related keywords (e.g., "wait", "alternatively"), underscoring the need for more reliable indicators of reasoning quality.
We next examine teacher distillation. A teacher model’s responses convey two main elements: (1) reasoning patterns and (2) domain knowledge. To disentangle their effects, we compare three models: the base model, the publicly released DeepSeek reasoning model, which is distilled on 800k responses from DeepSeek-R1 and likely incorporates substantial new knowledge, and our own reasoning-only model, trained only on teacher responses for questions where the base model is already able to produce correct answers. We find that both distilled models yield substantial accuracy gains, but only the DeepSeek model shows clear capability improvement. These results indicate that teacher distillation does not always expand capability, even when accuracy meaningfully improves. While further investigation is needed to confirm, we conjecture that this difference stems from whether new knowledge is introduced during distillation: introducing new knowledge may expand capability, whereas distilling only reasoning patterns improves accuracy but not capability. Interestingly, for the reasoning-only model, we also find that accuracy of the easier questions improves to the detriment of the most difficult questions, mirroring the RLVR.
Taken together, our findings provide a clearer picture of different dynamics in the model behavior during RLVR training and distillation, and offer insights into strategies for enhancing the fundamental abilities of LLMs.
@@ 2 Related Work
Training reasoning models. RLVR has emerged as a key method for training LLMs to tackle complex reasoning tasks by generating long CoTs DeepSeek-AI (2025); Lambert et al. (2024); OpenAI (2024). It has shown strong performance across model sizes Gandhi et al. (2025); Hu et al. (2025); Liu et al. (2025b); Xu et al. (2025); Yeo et al. (2025); Zeng et al. (2025) and domains Pan et al. (2025); Shrestha et al. (2025); Xie et al. (2025); Zhang et al. (2025). Numerous RLVR variants have also been proposed to improve performance, data efficiency, and computational cost Fatemi et al. (2025); Liu et al. (2025b); Shao et al. (2025, 2024); Wang et al. (2025a, b); Yu et al. (2025); Zuo et al. (2025). Distilling high-quality CoT data is another effective approach for enhancing LLM reasoning. Such data are obtained either by prompting large models Yu et al. (2024); Zelikman et al. (2022) or by human annotation of complex reasoning traces Qin et al. (2024); Xiang et al. (2025); Ye et al. (2025). A widely used strategy now involves distilling long CoT responses from RLVR-trained models into student models, often yielding substantial performance gains Huang et al. (2024); Min et al. (2024); Muennighoff et al. (2025); Shrestha et al. (2025). Our work examines both RLVR and distillation, and evaluates how these two approaches differentially shape reasoning behavior in LLMs.
Capability expansion in RLVR. There is ongoing debate about whether RLVR develops genuinely new capabilities not already present in a model. Several works Dang et al. (2025); Yue et al. (2025); Zhao et al. (2025) argue that RLVR merely amplifies correct reasoning already latent in the model. By contrast, ProRL demonstrates empirically that, given sufficient compute and diverse data, RLVR can enable models to solve previously unsolvable tasks in some domains-such as logic puzzles—suggesting the possibility of true capability expansion Liu et al. (2025a). OMEGA provides a more controlled analysis by carefully adjusting the knowledge and difficulty requirements of training and test math problems. Their results show that models can generalize to higher difficulty levels when the required knowledge is the same, but remain weak at chaining compositional skills or adopting novel, unconventional strategies Sun et al. (2025). Similarly, e3 finds that only problems with a sufficiently large verification–generation gap benefit from test-time scaling, through experimenting under settings where the problem types of training and test sets are strictly controlled Setlur et al. (2025). However, outside such carefully constrained conditions—in typical scenarios where both training and test sets consist of heterogeneous problems with uncontrolled knowledge and difficulty—studies consistently find that RLVR does not substantially expand capability. Theoretical analysis conducted by Wu et al. further argues that, in general, the shrinkage of empirical support outweighs its expansion in such scenarios, leading to little capability gain in RLVR Wu et al. (2025). In this work, we analyze RLVR under such general, uncontrolled math problem settings. By examining how accuracy shifts across difficulty levels, we show that RLVR often fails to improve capability as it tends to deliver gains on easier problems at the expense of performance on harder ones.
Reasoning pattern and knowledge in distillation. Several studies have examined the respective roles of domain knowledge and reasoning patterns in improving accuracy through distillation. For instance, Shrestha et al. distill teacher responses from logic puzzles—where domain knowledge is minimal—and show that transferring reasoning patterns alone can yield substantial performance gains across domains such as mathematics and coding Shrestha et al. (2025). Similarly, Huan et al. demonstrate that distilling math problem responses leads to notable improvements in reasoning tasks in other domains Huan et al. (2025). However, work on capability remains limited. Yue et al. suggest that distillation can drive capability expansion, but their analysis does not disentangle the effects of reasoning patterns and knowledge injection Yue et al. (2025). In contrast, our study explicitly controls for this distinction and investigates how each factor differentially influences model capability.
@@ 3.1 Formal Definitions
We evaluate models along two dimensions: accuracy and capability. Informally, accuracy measures how likely a model is to generate a correct answer in a single attempt, while capability measures whether a correct answer exists within the model’s response distribution.
@@ 4 Why Doesn’t RLVR Improve Capability?
Prior work has shown across multiple models that RLVR yields substantial gains in accuracy but often fails to improve capability, as measured by pass@$k$ with sufficiently large $k$ Yue et al. (2025). In this section, we extend this observation and analyze the phenomenon in greater depth. Specifically, we aim to answer: Why does RLVR raise accuracy while leaving capability unchanged or even degraded?
@@ 4.2 A Deeper-Dive: Analysis Based on Question Difficulty
To better understand the accuracy–capability dynamics of RLVR, we conduct a fine-grained analysis at the question difficulty level. For each question in the training and test sets, we generate 256 responses from the base model and compute its per-question success rate. Questions are then grouped into bins according to these rates: [0], [1–4], [5–16], [17–64], [65–128], and [129–256]. Within each bin, we collect the corresponding questions, retrieve the RL model’s responses to the same questions, and compute average success rates for both models. We then calculate the average success rate of the base and RL models in each bin and plot their absolute difference. The results are shown in Figure 1.
@@
We observe two consistent patterns. (1) For questions where the base model already has a moderately high success rate, the RL model shows substantial improvement. For example, in the training set, the [65–128] bin shows a 36.6% gain in average success rate relative to 256 responses, and the [17–64] bin shows a 24.7% gain. (2) In contrast, for questions where the base model has zero or near-zero success rate, the improvement is negligible. For instance, the [0] bin in the training set improves by only 0.5% points, and the [1–4] bin by just 0.6% points.
@@
Notably, we observe two clear trends. (1) Questions already in high-success bins tend to stay there or shift upward after RLVR. For example, in the [65–128] bin, 15.4% remain in place while 83.3% move to the top [129–256] bin; only 1.3% (1 question out of 78) drop lower. A similar upward shift appears in the [17–64] bin. (2) In contrast, questions in low-success bins—especially those near zero—tend to stagnate or regress. In the [1–4] bin, 61.1% remain and 16.7% fall to [0]; likewise, in the [5–16] bin, 44.2% stay or drop lower. This pattern shows a clearer picture of how RLVR fails to help previously unsolved questions and can even increase their number, as many with a small chance of being answered correctly end up never being solved after training.
To understand this behavior, we need to examine the internal dynamics of RLVR training. Take GRPO algorithm as an example. At each iteration, the model generates multiple responses (e.g., 8) per question in a batch. Each response is evaluated for correctness, and parameter updates are applied accordingly. If all 8 responses for a given question are incorrect, that question has no influence on the update. In contrast, for questions with a mix of correct and incorrect responses, the model parameters are nudged to increase the likelihood of generating correct answers and decrease incorrect ones. Now consider an early training step where the batch includes one difficult question (e.g., in bin [1–4]) and several easier ones (e.g., in bin [65–128]). With high probability, the model will generate only incorrect responses for the difficult question, while producing at least some correct answers for the easier ones. As a result, the parameter update will be guided entirely by the easier questions only. This can lead to the model assigning even lower probability to the correct answers for the difficult question, especially if model capacity is limited. This selective reinforcement continues throughout training, explaining why questions that initially had a small chance of being answered correctly may become even less likely to be solved after training. To confirm this selective reinforcement effect, we also analyze the entropy of the model’s output distributions—measured over 256 responses per question—across difficulty levels. We observe that entropy consistently decreases after RLVR training. Full results are provided in Appendix A.4.
To summarize, these results suggest the following insight: RLVR improves accuracy but not capability because RLVR focuses on improving the accuracy of the less-difficult questions to the detriment of the accuracy of the most difficult questions.
@@ 5.1 Self-Distillation Experiments
Table 1: Self-distillation results for the Qwen2.5-1.5B-Math base model and its RLVR-trained variant. The notation AA $\leftarrow$ $B$ indicates the student model $A$ is trained on responses from the teacher model $B$. Values in parentheses show gains over the student model with distillation.
@@
We take a step further and conduct another self-distillation experiment. Prior work has shown that distilling quality responses from a teacher model can effectively improve the performance of student model  Huang et al. (2024); Min et al. (2024); Muennighoff et al. (2025). In other words, responses that lead to large accuracy gains through distillation can be seen as having high quality  Kim et al. (2025). Based on this idea, we perform an additional experiment: distilling the RL model’s responses into the base model.
@@
In summary, when distilling the base model with correct responses from the base model there is only minor improvement, whereas when distilling with correct responses from the RLVR-trained model there is significant improvement, on par with RLVR itself. This suggests that there is a qualitative difference in the two response types, and reveals the following insight: RLVR does more than merely increase the success probability for the easier questions—it enables the model to produce quality responses that were not present in its output distribution before training.